# Hierarchical Mixture of Experts: Generalizable Learning for High-Level Synthesis

### v2

@@ Methods
$\boldsymbol{h_{i}^{\prime}}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{block}}\boldsymbol{h_{i}})_{j}\cdot\boldsymbol{W_{j}}\boldsymbol{h_{i}}\ \ (v_{i}\in V_{B}),$ (2)
@@
Models. Based on our pre-explorations, we use 4 experts in the low-level MoEs, and we set the regularization terms’ weights of both low-level and high-level MoEs to 5e-3. Our baselines include the SOTA GNN model, HARP (Sohrabizadeh et al. 2023); HARP+MAML (Bai et al. 2022), which applies meta-learning method MAML (Finn, Abbeel, and Levine 2017) on this task inspired by (Bai et al. 2022) to learn a more generalizable initialization of parameters; and ProgSG (Qin et al. 2024), which combines HARP and language model. Although hierarchical MoE brings many more parameters, it needs about 32 GB of memory and can still fit into a single GPU. HARP and our model contain 359,370 and 1,329,403 parameters respectively.
Evaluation. We train the regression and classification models on the source kernels. We want to mimic the domain transfer situation of having scarce but representative labeled data on the target kernels, so we use 50 data points per kernel to fine-tune our regression model, and roughly the same ratio of data points, 265 samples per kernel, to fine-tune the classification model. To select representative data points, we use K-means based on the graph representation. We conduct both offline and online evaluations. In offline evaluation, we calculate the fine-tuned regression model’s mean squared error (MSE) on the left-out data points in the target kernels for each regression objective and sum up the five objectives’ MSE. In the online evaluation, we use the DFS search used in previous studies (Sohrabizadeh et al. 2022, 2023) to search for pragma designs. We limit our search range to 75,000 pragma designs, since it typically takes an hour. We use the fine-tuned classification model to predict the validity and the fine-tuned regression model to predict the latency and resource utilization. From designs that are predicted to be valid and satisfy resource constraints, we choose the top 10 designs with the lowest predicted latencies to run HLS. We report the best design from the selected top-10 and the training dataset of the target kernels during fine-tune, since these are all labeled data points, and we calculate its speedup compared to the best design in the dataset. We run the offline evaluation five times and the online evaluation three times, then we report the mean results.
@@
Table 1 shows the main results. In the main experiments, we use the second design of the high-level gating network as it performs better. MAML could not perform well in this setting. It might be because we have many more source kernels compared to the previous paper that utilizes MAML on this task (Bai et al. 2022). Different kernels might result in different directions of the meta gradient, leading to unstable meta-learning. ProgSG which combines GNN and language models has a strong ability on the HLSyn dataset when the training data is sufficient (Qin et al. 2024), but it severely overfits in our data-scarce setting. During fine-tuning, the training loss is lower than 0.05, but the test loss is high. Comparatively, the hierarchical MoE is a more generalizable structure. In the online evaluation, on most kernels, the best of three single MoE models outperforms HARP. However, different kernels favor different low-level MoEs. For example, the block MoE model performs better than node MoE and graph MoE for “Fd”, while graph MoE is the best for “Gemv”, “Ja”, and “Tr”. By aggregating them together, the hierarchical MoE performs the best or close to the best on almost every kernel. Its geometric mean speedup achieves about 26.6% higher than HARP.
Table 3: Offline evaluation results on more complex kernels.
@@
Table 4: Using MoE on various granularities in a single model. “N”, “B”, and “G” represent node, block, and graph. Apart from selecting 50 representative points by K-means for fine-tuning, we also experiment with random data split. The hierarchical MoE model shows the highest stability when the data quality is worse due to random split.
@@
Table 5: Using MoE on a single granularity in the hierarchical MoE. “Node/Block/Graph” means only using MoE on the node/block/graph granularity in the hierarchical MoE. “All” means aggregating the three granularities.
@@
Table 6: Ablation study of two-stage training and high-level gating network. “Input” means that the high-level gating network utilizes the input features, and “Hidden” means that it utilizes the hidden representations.
@@
There are more ablation studies in the appendix. They verify that the hierarchical MoE’s performance is not due to increased parameter size or increased expert number, but the hierarchical structure and combination of three granularities. An ablation study also verifies that the current complicated design of the block MoE is better than some other variants.
@@
Table 7: Average assigned weights of the high-level MoE.
@@
We pick the best hierarchical MoE model based on the previous five repeated offline evaluation experiments to do a case study on the low-level MoE. Due to the limited space, here we only analyze the block MoE, and we analyze the node and graph MoEs in the appendix. The pragmas modify the pseudo nodes, so block MoE is a window for us to analyze the pragmas. We summarize the weights of each expert for each pragma type in Table 8. There are three types of pragmas: loop tiling, pipeline, and parallelization. The third expert is good at dealing with loop tiling, pipeline, and small parallel factors; the second expert is good at dealing with large parallel factors; the other two experts deal more with pseudo nodes that do not have pragmas. Different experts diversify their roles, which improves the generalizability.
@@ Conclusion
Domain generalization is a big challenge for HLS prediction models. Based on the unique challenges and opportunities, we propose the hierarchical MoE structure, where different expert networks can specialize in different regions, and it can be flexibly adapted to any GNN. In the low-level MoE, we apply MoE on the three natural granularities of the graph: node, basic block, and graph. In the high-level MoE, we aggregate the three low-level MoE models, so that different data points can flexibly decide which one to use. To address the severe expert polarization, we propose a two-stage training strategy. Extensive experiments have verified its effectiveness. Nonetheless, the generalizability of HLS prediction models still remains a big challenge, and future works can further improve the performance.
@@ Appendix B Additional Problem Analysis
Second, the labeling is time-consuming on the new kernels. Each run takes minutes to hours (Sohrabizadeh et al. 2023). Therefore, we use the constraint of data scarcity of only using 50 data points per kernel to finetune the model on new kernels.
@@ Appendix D Supplementary Experiment Results
Table 13: Influence of the number of experts in the single-level MoE model. The listed results are test MSE. We use bold to mark the best number of experts for each model.

### v3

@@ Methods
$\boldsymbol{h_{i}^{\prime}}=\sum_{j=1}^{n}softmax(\boldsymbol{W_{G}^{block}}\boldsymbol{h_{i}})_{j}\cdot\boldsymbol{W_{j}}\boldsymbol{h_{i}}\ \ (v_{i}\in V_{B}),\vspace{-1mm}$ (2)
@@
Models. Based on our pre-explorations, we use 4 experts in the low-level MoEs, and we set the regularization terms’ weights of both low-level and high-level MoEs to 5e-3. Our baselines include the SOTA GNN model, HARP (Sohrabizadeh et al. 2023); HARP+MAML (Bai et al. 2022), which applies meta-learning method MAML (Finn, Abbeel, and Levine 2017) on HARP inspired by (Bai et al. 2022) to learn a more generalizable initialization of parameters; and ProgSG (Qin et al. 2024), which combines HARP and language model. Although hierarchical MoE brings many more parameters, it needs about 32 GB of memory and can still fit into a single GPU. HARP and our model contain 359,370 and 1,329,403 parameters respectively.
Evaluation. We train the regression and classification models on the source kernels. We want to mimic the domain transfer situation of having scarce but representative labeled data on the target kernels, so we use 50 data points per kernel to fine-tune our regression model, and roughly the same ratio of data points, 265 samples per kernel, to fine-tune the classification model. To select representative data points, we use K-means based on the graph representation. We conduct both offline and online evaluations. In offline evaluation, we calculate the fine-tuned regression model’s mean squared error (MSE) on the left-out data points in the target kernels for each regression objective and sum up the five objectives’ MSE. In the online evaluation, we use the DFS search used in previous studies (Sohrabizadeh et al. 2022, 2023) to search for pragma designs. We limit our search range to 75,000 pragma designs, since it typically takes an hour. We use the fine-tuned classification model to predict the validity and the fine-tuned regression model to predict the latency and resource utilization. From designs that are predicted to be valid and satisfy resource constraints, we choose the top 10 designs with the lowest predicted latencies to run HLS. We report the best design from the selected top-10 and the training dataset of the target kernels during fine-tuning, since these are all labeled data points, and we calculate its speedup compared to the best design in the dataset. We run the offline evaluation five times and the online evaluation three times, then we report the mean results.
@@
Table 1 shows the main results. In the main experiments, we use the second design of the high-level gating network as it performs better. MAML could not perform well in this setting. It might be because we have many more source kernels compared to the paper that proposes to use MAML on this task (Bai et al. 2022). Different kernels might result in different directions of the meta gradient, leading to unstable training. ProgSG which combines GNN and language models has a strong ability when the training data is sufficient, but it overfits in the data-scarce setting. During fine-tuning, the training loss is lower than 0.05, but the test loss is high. Comparatively, the hierarchical MoE is more generalizable. In the online evaluation, on most kernels, the best of three single MoE models outperforms HARP. However, different kernels favor different low-level MoEs. For example, the block is the best granularity for “Fd”, while the graph is the best granularity for “Gemv”, “Ja”, and “Tr”. By aggregating them together, hierarchical MoE performs the best or close to the best on almost every kernel.
@@
Table 3: Offline evaluation results on more complex kernels.
@@
Table 4: Using MoE on various granularities in a single model. “N”, “B”, and “G” represent node, block, and graph.
@@
Table 5: Using MoE on a single granularity in the hierarchical MoE. “Node/Block/Graph” means only using MoE on the node/block/graph granularity in the hierarchical MoE.
@@
Table 6: Ablation study of two-stage training and high-level gating network. “Input” is the first design of the high-level gating network, while “Hidden” is the second design.
@@
More ablation studies are in the appendix. They verify that the hierarchical MoE’s performance is not due to increased parameter size or expert number, but the hierarchical structure and combination of three granularities.
@@
Table 7: Average assigned weights of the high-level MoE.
@@
Table 8: Average assigned weights of the block MoE. Each column shows the expert weights for pseudo nodes modified by a certain pragma type.
We pick the best hierarchical MoE model based on the previous five repeated offline evaluation experiments to do a case study on the low-level MoE. Due to the limited space, here we only analyze the block MoE, and we analyze the node and graph MoEs in the appendix. The pragmas modify the pseudo nodes, so block MoE is a window for us to analyze the pragmas. We summarize the weights of each expert for each pragma type in Table 8. There are three types of pragmas: loop tiling, pipeline, and parallelization. The third expert is good at dealing with loop tiling, pipeline, and small parallel factors; the second expert is good at dealing with large parallel factors; the other two experts deal more with pseudo nodes that do not have pragmas. Different experts diversify their roles, which improves generalizability.
@@ Conclusion
Domain generalization is a big challenge for HLS prediction models. Based on the unique challenges and opportunities, we propose the hierarchical MoE structure. In the low-level MoE, we apply MoE on one of the three natural granularities of the graph: node, basic block, or graph. In the high-level MoE, we aggregate the three low-level MoE models, so that different data points can flexibly decide which one to use. To address the severe expert polarization, we propose a two-stage training strategy. Extensive experiments have verified its effectiveness. Nonetheless, the generalizability of HLS prediction models still remains a big challenge, and future works can further improve the performance.
@@ Appendix B Additional Problem Analysis
Second, labeling the new kernel’s data points by HLS is time-consuming. Each run takes minutes to hours (Sohrabizadeh et al. 2023). Therefore, we use the constraint of data scarcity of only using 50 data points per kernel to finetune the model on new kernels.
@@ Appendix D Supplementary Experiment Results
Table 13: Influence of the number of experts in the single-level MoE model. The listed results are test MSE.